The sum of two consecutive even integers is 86. Find the larger number. I don't really know how to do this.
Answer by Alan3354(69443)

The sum of two consecutive even integers is 86. Find the larger number. I don't really know how to do this.
--------------
n + n+2 = 86
2n+2 = 86
2n = 84
n = 42
42 + 44 = 86